# GaussianAnything: Interactive Point Cloud Latent Diffusion for 3D Generation

### v1

@@ 4 GaussianAnything
    A conditional cascaded transformer denoiser $\bm{\epsilon}_{\Theta}^{h}({\mathbf{z}}_{h,t},{\mathbf{z}}_{x,0},t,c)\circ\bm{\epsilon}_{\Theta}^{x}({\mathbf{z}}_{x,t},t,c)$ to denoise the noisy latent code ${\mathbf{z}}_{t}$ given diffusion time step $t$ and condition prompt $c$;
@@
Beyond the advantages shared by existing 3D LDM (Zhang et al., 2024; Lan et al., 2024), our design offers a flexible point-cloud structured latent space and enables interactive 3D editing.
In the following subsections, we first discuss the proposed 3D VAE with a detailed framework design in Sec 4.1. Based on that, we introduce the cascaded conditional 3D diffusion stage in Sec. 4.2. The method overview is shown in Fig. 2.
4.1 Point-cloud Structured 3D VAE
Unlike image and video, the 3D domain is un-uniform and represented differently for different purposes. Thus, how to encode 3D objects into the latent space for diffusion learning plays a crucial role in the 3D generation performance. This challenge is two-fold: what 3D representations to encode, and what network architecture to process the input.
Versatile 3D Input. Instead of using dense point cloud (Zhang et al., 2024; Li et al., 2024), we adopt multi-view posed RGB-D(epth)-N(ormal) images as input, which encode the 3D input more comprehensively and can be efficiently processed by well-established network architectures (Sajjadi et al., 2022; Wu et al., 2023) in a flexible manner. Specifically, the input is a set of multi-view renderings $\mathcal{R}$ of a 3D object, where each rendering within the set $R=(I,0pt,{N},\pi)$ contains thorough 3D attributes that depict the underlying 3D object from the given viewpoint: the rendered RGB image $I\in^{H\times W\times 3}$, depth map $0pt\in^{H\times W}$, normal map ${N}\in^{H\times W\times 3}$, and the corresponding camera pose $\pi$.
To unify these 3D attributes in the same format, we further process the camera $\pi$ into Plücker coordinates (Sitzmann et al., 2021) ${\mathbf{p}}_{i}=({\mathbf{o}}\times{\mathbf{d}}_{u,v},{\mathbf{d}}_{u,v})\in^{6}$, where ${\mathbf{o}}_{i}\in^{3}$ is the camera origin, ${\mathbf{d}}_{u,v}\in^{3}$ is the normalized ray direction, and $\times$ denotes the cross product. Thus, the Plücker embedding of a given camera $\pi$ can be expressed as ${\mathbf{P}}\in^{H\times W\times 6}$. Besides, following MCC (Wu et al., 2023), we use $\pi$ to unproject the depth map into their 3D positions $X\in^{H\times W\times 3}$. The resulting information is channel-wise concatenated, giving $\tilde{R}=[I\oplus X\oplus N\oplus{\mathbf{P}}]\in^{H\times W\times(3+3+3+6=15)}$.
Transformer-based 3D Encoding. Given the 3D renderings $\mathcal{R}$, encoding them into a 3D latent space remains a significant challenge. Independently processing each input rendering $\tilde{R}$ with existing network architecture (Wu et al., 2023; Dosovitskiy et al., 2021) overlooks the information from other views, leading to 3D inconsistency and content drift across views (Liu et al., 2023b).
${\mathbf{z}}_{z}=\mathcal{E}_{\bm{\phi}}^{\text{TX}}(\mathcal{E}_{\bm{\phi}}^{\text{CNN}}(\{\tilde{R})\}),$ (5)
where ${\mathbf{z}}_{z}$ is the set latent corresponding to the 3D input. This can be seen as the full-attention version of existing 3D attention-augmented architecture. The resulting latent codes ${\mathbf{z}}_{z}$ fully capture the intact 3D information corresponding to the input. Compared to existing work that adopts point clouds only as input (Zhang et al., 2024; Li et al., 2024), our proposed solution supports more 3D properties as input in a flexible way. Besides, the attention operations can be well optimized over modern GPU architecture (Dao et al., 2022; Dao, 2024).
Point Cloud-structured Latent Space. Though ${\mathbf{z}}_{z}$ fully captures the given 3D input, it cannot be directly used for diffusion learning due to the lack of a regularized, low-rank latent space (Rombach et al., 2022). Specifically, ${\mathbf{z}}_{z}$ has a shape of $V\times(H/f)\times(W/f)\times C$, where $V$ is the number of views used for input, $H,W$ is the input resolution and $f$ is the down-sampling factor of the CNN backbone. Given $V=8$, $f=8$, and $H=W=512$, the resulting latent codes will have a shape of $32768\times C$, which are unsuitable for diffusion learning. Moreover, using the set (Lee et al., 2019) as the 3D latent space (Zhang et al., 2023a; 2024; Li et al., 2024) also sacrifices an explicit, editable latent space (Mou et al., 2023a) for flexibility.
Here, we resolve these issues by proposing a point cloud-structured latent space. Specifically, we project the un-structured features ${\mathbf{z}}_{z}$ onto the manifold of the input 3D shape through the cross attention layer:
where $\text{CrossAttn}(Q,K,V)$ denotes a cross attention block with query $Q$, key $K$, and value $V$. ${\mathbf{z}}_{x}\in^{3\times N}$ is a sparse point cloud sampled from the surface of the 3D input with Farthest Point Sampling (FPS) (Qi et al., 2017), and PE denotes positional embedding (Tancik et al., 2020). Intuitively, we define a read cross attention block (Huang et al., 2024b) that cross attends information from un-structured representation ${\mathbf{z}}_{z}$ into the point-cloud structured feature ${\mathbf{z}}_{h}\in^{C_{h}\times N}$, with $C_{h}\ll C$. In this way, we obtain the point-cloud structured latent code ${\mathbf{z}}=[{\mathbf{z}}_{x}\oplus{\mathbf{z}}_{h}]\in^{(3+C_{h})\times N}$ for diffusion learning.
High-quality 3D Gaussian Decoding. Given the point cloud-structured latent codes, how to decode them into high-quality 3D representation for supervision remains challenging. Though dense point cloud (Huang et al., 2024b) is a straightforward solution, it fails to depict high-quality 3D structure with limited point quantity. Here, we resort to surfel Gaussian (Huang et al., 2024a), an augmented point-based 3D representation that supports high-fidelity 3D surface modeling and efficient rendering. Specifically, our decoder first decodes the input through the 3D-DiT blocks (Peebles & Xie, 2023; Lan et al., 2024), which has shown superior performance against traditional transformer layer:
$\tilde{{\mathbf{z}}}\coloneq\mathcal{D}_{T}(\text{MLP}({\mathbf{z}})),$ (7)
To achieve denser Gaussians prediction, we cascade the upsampling transformer defined in Eq. (8) for $K$ times, giving the final Upsampler $\mathcal{D}_{U}$ for high-quality Gaussian output. Note that our solution outputs a set of Gaussians that are uniformly distributed on the 3D object surface with near $100\%$ Gaussian utilization ratio. Existing pixel-aligned Gaussian prediction models (Tang et al., 2024; Yinghao et al., 2024; Szymanowicz et al., 2023), however, usually waste $50\%$ Gaussians due to view overlaps and empty background color. Besides, our intermediate Gaussians output naturally serves as $K$ LoD (Takikawa et al., 2021), which can be used in different scenarios to balance the rendering speed and quality.
where $\mathcal{L}_{\text{render}}$ is a mixture of the $L_{1}$ and VGG loss (Zhang et al., 2018), $\mathcal{L}_{\text{geo}}$ improves geometry reconstruction (Huang et al., 2024a), $\mathcal{L}_{\text{KL}}$ is the KL-reg loss (Kingma & Welling, 2013; Rombach et al., 2022) to regularize a structured latent space, and $\mathcal{L}_{\text{GAN}}$ improves perceptual fidelity. All loss terms except $\mathcal{L}_{\text{KL}}$ are applied over a randomly chosen LoD in each iteration, and the $\mathcal{L}_{\text{render}}$ is applied to both input-view and randomly sampled novel-view images. Besides, the geometry loss $\mathcal{L}_{\text{geo}}$ is composed of two regularization terms, including the depth distortion loss to concentrate the weight distribution along rays, inspired by Mip-NeRF (Barron et al., 2021; 2022). Given a ray of pixel, the distortion loss is defined as
$\mathcal{L}_{d}=\sum_{i,j}\omega_{i}\omega_{j}|d_{i}-d_{j}|,$ (10)
$\mathcal{L}_{n}=\sum_{i}\omega_{i}(1-\hat{{N}}_{i}^{\mathrm{T}}{N}),$ (11)
4.2 Cascaded 3D Generation with Flow Matching
After training the point-cloud structured 3D VAE, we get a dataset of $D$ shapes paired with condition vectors (e.g., caption or images), $\{({\mathbf{z}}_{i},c_{i})\}_{i\in[D]}$, where the shape is represented by latent code ${\mathbf{z}}$ through the 3D VAE aforementioned. Our goal is to train a flow-matching generative model to learn a diffusion prior on top of it. Below we present how we adapt flow-based models to our case.
Cascaded Flow Matching over Symmetric Data. As detailed in Sec. 3, flow matching involves training a neural network $\bm{\epsilon}_{\Theta}$ to predict the velocity $v$ of the noisy input ${\mathbf{z}}_{t}$ with the straight-line trajectory. After training, $\bm{\epsilon}_{\Theta}$ can sample from a standard Normal prior ${\mathcal{N}}(0,I)$ by solving the reverse ODE/SDE (Karras et al., 2022). In our case, the training data point is the point-cloud structured latent code ${\mathbf{z}}=[{\mathbf{z}}_{x}\oplus{\mathbf{z}}_{h}]\in^{(3+C_{h})\times N}$, which is symmetric and permutation invariant (Zeng et al., 2022; Nichol et al., 2022). Based on this property, we opt for diffusion transformer (Peebles & Xie, 2023) without positional encoding as the $\bm{\epsilon}_{\Theta}$ parameterization.
Here, rather than modeling ${\mathbf{z}}_{x}$ and ${\mathbf{z}}_{h}$ jointly, we empirically found that a cascaded framework (Ho et al., 2021; Lyu et al., 2024; 2023) leads to better performance. Specifically, a conditioned sparse point cloud generative model $\bm{\epsilon}_{\Theta}^{x}$ is first trained to generate the overall structure of the given object:
The detailed cascading process is detailed Fig. 3 (c). Our proposed design enables better geometry-texture disentanglement and facilitates 3D editing over specific shape properties.
Conditioning Mechanism. Compared to LRM (Hong et al., 2024b; Tang et al., 2024) line of work which only handles image conditions, our native diffusion-based method enables more flexible 3D generation from diverse conditions. As shown in Fig. 3 (a-b), for the text-conditioned model, we adopt CLIP (Radford et al., 2021) to extract penultimate tokens as the condition embeddings; and for the image conditioned model, we use DINOv2 (Oquab et al., 2023) to extract global and patch features. All conditions are injected into the DiT architecture through a pre-norm (Xiong et al., 2020) cross-attention block. All the models are trained with Classifier-free Guidance (CFG) (Ho, 2021) by randomly dropping the conditions with a probability of $10\%$.
To cascade two diffusion models, we encode the output of stage-1 model $\bm{\epsilon}_{\Theta}^{x}$ with $\text{PE}({\mathbf{z}}_{x})$ as in Eq. (6), and add it to the first-layer features of $\bm{\epsilon}_{\Theta}^{h}$. This guarantees that generated features are paired with the input sparse point cloud structure.
5 Experiments
Datasets. To train our 3D VAE, we use the renderings provided by G-Objaverse (Qiu et al., 2023; Deitke et al., 2023) and choose a high-quality subset with around $176K$ 3D instances, where each consists of $40$ random views with RGB, normal, depth map and camera pose. For text-conditioned diffusion training, we use the caption provided by Cap3D (Luo et al., 2023; 2024) and 3DTopia Hong et al. (2024a). For image-conditioned training, we randomly select an image in the dataset of the corresponding 3D instance as the condition.
Implementation Details. For 3D VAE, we choose $V=8$ views of RGB-D-N renderings as input to guarantee a thorough coverage of the 3D object. The CNN Encoder is implemented with a similar architecture as LDM VAE (Rombach et al., 2022) with a down-sampling factor of $f=8$, and the multi-view transformer has five layers as in RUST (Sajjadi et al., 2023). The sparse point cloud ${\mathbf{z}}_{x}$ has a size of $N\times 3$ where $N=768$, and the corresponding featured point cloud ${\mathbf{z}}_{h}$ has a dimension of $N\times 10$. For upsampling blocks, we employ $K=3$ blocks with $f_{u}^{1}=8$, $f_{u}^{1}=4$, and $f_{u}^{1}=3$, giving $73,768$ Gaussians in total. All transformer blocks follow a pre-norm design (Xiong et al., 2020). During 3D VAE training, the model is supervised by randomly chosen LoD renderings, with $\lambda_{\text{kl}}=2e-6$, $\lambda_{\text{d}}=1000$, $\lambda_{\text{n}}=0.2$, and $\lambda_{\text{GAN}}=0.1$. We adopt batch size $64$ with both input and random novel views for training. During the conditional diffusion training stage, we adopt batch size $256$. All models are efficiently and stably trained with lr $=1e-4$ on 8×8$\times$8 ×A100 GPUs for $1M$ iterations with BF16 and FlashAttention (Dao, 2024) enabled. We use CFG=$4$ and $250$ ODE steps for all sampling results.
5.1 Metrics and Baselines
Evaluating Image-to-3D Generation. We evaluate GaussianAnything on both image and text conditioned generation. Regarding image-conditioned 3D generation methods, we compare the proposed method with three lines of methods: single-image to 3D methods: OpenLRM (He & Wang, 2023; Hong et al., 2024b), Splatter Image (Szymanowicz et al., 2023), multi-view images to 3D methods: One-2-3-45 Liu et al. (2023a), CRM (Wang et al., 2024), Lara (Chen et al., 2024), LGM (Tang et al., 2024), and native 3D diffusion models: Shape-E (Jun & Nichol, 2023) and LN3Diff-image (Lan et al., 2024).
Quantitatively, we benchmark rendering metrics with FID (Heusel et al., 2017), KID (Bińkowski et al., 2018), and MUSIQ (Ke et al., 2021; Zhou et al., 2022). For 3D quality metrics, we adopt Point cloud FID (P-FID), Point cloud KID (P-KID), Coverage Score (COV), and Minimum Matching Distance (MMD) as the metrics. Following previous works Nichol et al. (2022); Zhang et al. (2023a); Yariv et al. (2024), we adopt the pre-trained PointNet++ provided by Point-E (Nichol et al., 2022) for calculating P-FID and K-FID. Qualitatively, GSO (Downs et al., 2022; Zheng & Vedaldi, 2023) dataset is used for visually inspecting image-conditioned generation results.
Evaluating Text-to-2D Generation. Regarding text-conditioned 3D generation methods, we compare against Point-E (Nichol et al., 2022), Shape-E (Jun & Nichol, 2023), 3DTopia (Hong et al., 2024a), and LN3Diff-text (Lan et al., 2024). CLIP score (Radford et al., 2021) is reported following the previous works (Lan et al., 2024; Hong et al., 2024a).
Table 1: Quantitative evaluation of image-conditioned 3D generation. Here, quality of both 2D rendering and 3D shapes is evaluated. As shown below, the proposed method demonstrates strong performance across all metrics. Although multi-view images-to-3D approaches like LGM achieves better performance on the FID/KID metrics, they fall short on more advanced image quality assessment metrics such as MUSIQ and performs significantly worse in 3D shape quality. For multi-view to 3D methods, we also include the number of input views (V=##\##).
[…]
OpenLRM 38.41 1.87 45.46 35.74 12.60 39.33 29.08
Splatter-Image 48.80 3.65 30.33 19.72 7.03 37.66 30.69
One-2-3-45 (V=12) 88.39 6.34 59.02 72.40 30.83 33.33 35.09
CRM (V=6) 45.53 1.93 64.10 35.21 13.19 38.83 28.91
Lara (V=4) 43.74 1.95 39.37 32.37 12.44 39.33 28.84
LGM (V=4) 19.93 0.55 54.78 40.17 19.45 50.83 22.06
Shape-E 138.53 11.95 31.51 20.98 7.41 61.33 19.17
LN3Diff 29.08 0.89 50.39 27.17 10.02 55.17 19.94
Ours 24.21 0.76 65.17 8.72 3.22 59.50 15.48
5.2 Evaluation
Image-to-3D Generation. Our proposed framework enables 3D generation given single-view image conditions, leveraging the architecture detailed in Fig. 3 (b). Following current method (Chen et al., 2024; Tang et al., 2024), we qualitatively benchmark our method in Fig. 4 over the single-view 3D reconstruction task on the unseen images from the GSO dataset. Our proposed framework is robust to inputs with complicated structures (row 1-2), low-light (row 3), and self-occlusion (row 4), yielding consistently intact 3D reconstruction. Besides, our generative-based method shows a more natural back-view reconstruction, as opposed to regression-based methods that are commonly blurry on uncertain areas (row 5).
Quantitatively, we showcase the evaluation in Tab. 1. As can be seen, our proposed method achieves state-of-the-art performance over all 3D metrics, and shows competitive results over conventional 2D rendering metrics FID/KID. Note that LGM leverages pre-trained MVDream (Shi et al., 2023b) as the first-stage generation, and then maps the generated $4$ views to pixel-aligned 3D Gaussians. This cascaded pipeline achieves better visual quality, but prone to yield distorted 3D geometry, as visualized in Fig. 4.
Table 2: Quantitative Evaluation on Text-to-3D. The proposed method outperforms existing methods on CLIP scores over two different backbones.
Method ViT-B/32 ViT-L/14
Point-E 26.35 21.40
Shape-E 27.84 25.84
LN3Diff 29.12 27.80
3DTopia 30.10 28.11
Ours 31.80 29.38
Text-to-3D Generation. We demonstrate the text-to-3D generation performance in Fig. 5 and Tab. 2. The diffusion model trained on GaussianAnything’s latent space has demonstrated high-quality text-to-3D generation of generic 3D objects, yielding superior performance in terms of object structure, textures, and surface normals. Quantitatively, our proposed method achieves better text-3D alignment against competitive baselines.
5.3 Ablation Study and Analysis
[…]
3D VAE Design. In Tab. 5.3, we benchmark each component of our 3D VAE architecture over a subset of Objaverse with $50K$ instances and record the LPIPS at $100K$ iterations. As shown in Tab. 5.3, our input design performs better against dense ($16,384$) colored point cloud (Zhang et al., 2024), and the reconstruction quality consistently improves by including normal map as input and cascading more Gaussian upsampling blocks.
Figure 7: Qualitative ablation of Cascaded diffusion and latent space editing. We first show the effectiveness of our two-stage cascaded diffusion framework in (a). Compared to Fig. 5, the single-stage 3D diffusion yields worse texture details and 3D structure intactness. In (b), we validate the latent point cloud editing yields less 3D artifacts compared to direct 3D editing on the 3D Gaussians.
Effectiveness of Cascaded 3D Diffusion. We qualitatively ablate the cascaded 3D diffusion choice in Fig. 7 (a), where a single text-conditioned DiT is trained to synthesize the 3D point cloud and features jointly. Clearly, the jointly trained model has a worse texture with 3D shape artifacts to our cascaded design. Besides bringing better editing capability as shown in Fig. 6, our cascaded design enables more flexible training, where the models of two stages can be trained in parallel.
3D Editing on the 3D Latent Space. Finally, we ablate the 3D editing performance in Fig. 7 (b). As can be seen, direct editing on the final Gaussians leads to 3D artifacts, while editing on our 3D latent space yields more holistic and cleaner results.
6 Conclusion and Discussions
In this work, we present a new paradigm of 3D generative model by learning the diffusion model over a interactive 3D latent space. A dedicated 3D variational autoencoder encodes multi-view 3D attributes renderings into a point-cloud structured latent space, where multi-modal diffusion learning can be efficiently performed. Our framework achieves superior performance over both text- and image-conditioned 3D generation, and potentially facilitates numerous downstream applications in 3D vision and graphics tasks.
Limitations and Future Work. Our method comes with some limitations to be resolved. 3D VAE side, we observe that the reconstruction textures are sometimes blurry on 3D objects with intricate textures. A potential solution is to leverage pixel-aligned features (Saito et al., 2019; Lan et al., 2023a; Melas-Kyriazi et al., 2023) to alleviate this issue. Moreover, incorporating rendering loss (Anciukevičius et al., 2023) during diffusion training may further boost the generation quality. Besides, adding more real-world data such as MVImageNet Yu et al. (2023) and more control conditions Zhang et al. (2023b) is also worth exploring.
Appendix
Appendix A Implementation details
A.1 Training details
VAE Architecture. For the convolutional encoder $\mathcal{E}_{\bm{\phi}}$, we adopt a lighter version of LDM Rombach et al. (2022) encoder with channel $64$ and $1$ residual blocks for efficiency. When training on Objaverse with $V=8$, we incorporate 3D-aware attention Shi et al. (2023b) in the middle layer of the convolutional encoder. The multi-view transformer architecture is similar to RUST (Sajjadi et al., 2023; 2022). For each upsampler $\mathcal{D}_{U}^{k}$, we have $2$ transformer blocks in the middle. All hyper-parameters remain at their default settings. Regarding the transformer decoder $\mathcal{D}_{T}$, we employ the DiT-B/2 architecture due to VRAM constraints. Compared to LN3Diff (Lan et al., 2024), we do not adopt cross-plane attention in the transformer decoder.
A.2 Data and Baseline Comparison
Appendix B More visual results and videos
Please check our supplementary video demo and attached folders for more results.
Broader Social Impact. In this paper, we introduce a new latent 3D diffusion model designed to produce high-quality surfel Gaussians using a single model. As a result, our approach has the potential to be applied to generating DeepFakes or deceptive 3D assets, facilitating the creation of falsified images or videos. This raises concerns as individuals could exploit such technology with malicious intent, aiming to spread misinformation or tarnish reputations.

### v2

@@ Appendix B Discussions of limitations
    Enhancing the 3D VAE Quality.. The performance of the 3D VAE could be improved by increasing the number of latent points and incorporating a pixel-aligned 3D reconstruction paradigm, such as PiFU (Saito et al., 2019), to achieve finer-grained geometry and texture alignment.
@@
    Incorporating Additional Losses in Diffusion Training. Currently, the diffusion training relies solely on latent-space flow matching. Prior work, such as DMV3D (Xu et al., 2024c), demonstrates that incorporating a rendering loss can significantly enhance the synthesis of high-quality 3D textures. Adding reconstruction supervision during diffusion training is another promising avenue to improve output fidelity.
@@
    Leveraging 2D Pre-training Priors. At present, the models are trained exclusively on 3D datasets and do not utilize 2D pre-training priors as effectively as multi-view (MV)-based 3D generative models. A potential improvement is to incorporate 2D priors more effectively, for instance, by using multi-view synthesized images as conditioning during training instead of single-view images.
  4. 4.
    Expanding Dataset Diversity. Utilizing more diverse and extensive 3D datasets, such as Objaverse-XL (Deitke et al., 2023a) and MVImageNet (Yu et al., 2023), could further enhance the quality and generalizability of 3D generation.
  5. 5.
    Support of Physics-based Rendering. Currently, our proposed method leverages 2DGS as the underlying 3D representation since it balances the 2D rendering and 3D surface quality. However, incorporating more advanced 3D representations such as 3DGRT (Moenne-Loccoz et al., 2024) can further support physics-based ray tracing over 3D Gaussians. Besides, this can be achieved by directly fine-tuning our pre-trained 3D VAE decoder only with the PBR rendering pipeline.
By addressing these aspects, the proposed method could achieve significant advancements in both the quality and versatility of its 3D generation capabilities.
Appendix C Discussions of Concurrent Work
After the submission of this paper, several related works on 3D generation have been released. We discuss them here.
FreeSplatter (Xu et al., 2024b) extends the multi-view generation and 3D reconstruction pipeline for 3D object generation, eliminating the need for input poses. It also supports 3D scene generation by training on the scene-level datasets (Yao et al., 2020; Dai et al., 2017). While following a different paradigm, our method could benefit from its pose-free design to enable image-conditioned 3D generation from multiple input views.
Geometric Distribution (Zhang et al., 2024a) introduces a novel 3D representation by learning a point cloud diffusion model for single-instance objects. This approach closely resembles our stage-1 diffusion model but focuses on encoding geometric information. While it shows strong potential for downstream tasks, its efficiency remains a challenge for broader applications.
AtlasGaussians (Yang et al., 2025), like our method, proposes a native 3D diffusion model over 3D Gaussians. It employs point clouds as input for 3D VAE compression and utilizes local patch features for 3D Gaussian decoding. However, it lacks an explicit 3D latent space, making it unsuitable for interactive 3D editing.
Trellis (Xiang et al., 2024) adopts a similar cascaded 3D flow matching framework and achieves remarkable results in object generation. Unlike our method, which employs a point cloud latent space, Trellis leverages a sparse voxel structure with efficient operators for high-quality and fast 3D generation. It also supports interactive 3D editing via its explicit latent space. We hope this design, along with our proposed method, can establish a canonical paradigm for future 3D native generative models.
Appendix D More visual results and videos
Please also check our supplementary video demo and attached folders for video results.
Overview of the Qualitative Performance.
In Fig. 7, we include an overview of the qualitative performance of the proposed method, GaussianAnything. Here, we show the single-view conditioned 3D generation, text-conditioned 3D generation, and 3D-aware editing capabilities.
3D VAE Reconstruction.
Figure 8: 3D VAE Reconstruction. Here, we visualize the 3D VAE reconstruction performance across different level of details (LoD). As shown, higher LoD results in sharper textures and smoother surface. Better zoom in.
In Fig. 8, we include the 3D VAE reconstruction results of our model at $3$ level of details (LoD). Thanks to the versatile multi-view 3D attributes input and transformer design, our 3D VAE enables high-quality 3D reconstruction with visually attractive textures and smooth surface. The encoded point cloud-structured latent codes, ${\mathbf{z}}$, serves as a compact proxy for efficient 3D diffusion training. Besides, the 2D Gaussians Upsampler naturally facilitates LoD and facilitates speed / quality trade off in practice.
More Text-to-3D results.
Figure 9: More Qualitative Comparison of Text-to-3D. We present more text-conditioned 3D objects generated by GaussianAnything, alongside comparisons with competitive alternatives, including Point-E, Shape-E, and 3DTopia. As demonstrated, our approach consistently achieves superior quality in geometry, texture, and alignment between text and 3D content.
In Fig. 9, we present additional qualitative comparisons of text-to-3D generation with GaussianAnything. For this evaluation, we use relatively complex captions as input conditions and display two random samples generated by our model. As shown, GaussianAnything produces visually appealing results that characterized rich textures, smooth surface, and notable diversity. To further demonstrate the generality of our proposed method, in Fig. 10 we include the uncurated text-to-3D results over DF-415 (Poole et al., 2022) prompts with captions and more detailed descriptions.
Figure 10: More Qualitative Results of Text-to-3D over DF-415 Captions. Our proposed method generalizes to long captions with detailed descriptions. All results are uncurated.
Figure 11: Cascaded Text-to-3D with Point-E. Thanks to our cascaded 3D generation design, the stage-2 diffusion model in GaussianAnything seamlessly integrates with other point cloud generative models. To illustrate this capability, we leverage the state-of-the-art Point-E model. As demonstrated, our stage-2 diffusion model effectively transforms the point clouds generated by Point-E into diverse surfel Gaussians with more visually appealing features and enhanced geometric details.
Point-to-3D Generation with Cascaded Point-E. Moreover, the cascaded design of our stage-2 diffusion model, $\bm{\epsilon}_{\Theta}^{h}$, enables flexible 3D generation given point clouds from diverse sources. To demonstrate this capability, we integrate the output of a state-of-the-art 3D point cloud generative model, such as Point-E (Nichol et al., 2022), into the GaussianAnything generation pipeline. Specifically, we first generate the point cloud using Point-E based on a caption condition $c$. This generated point cloud is then used as input ${\mathbf{z}}_{x}$ to our stage-2 point cloud/text-conditioned diffusion model $\bm{\epsilon}_{\Theta}^{h}$. As shown in Fig. 11, the generated surfel Gaussians exhibit significantly improved texture quality and geometry fidelity compared to the Point-E point cloud outputs. This capability broadens the applicability of our method, enabling it to benefit from recent advances in point cloud generation (Huang et al., 2024b) and mesh generation (Siddiqui et al., 2023; Chen et al., 2024c) for producing high-quality object-level surfel Gaussians.
Broader Social Impact. In this paper, we introduce a new latent 3D diffusion model designed to produce high-quality surfel Gaussians using a single model. As a result, our approach has the potential to be applied to generate DeepFakes or deceptive 3D assets, facilitating the creation of falsified images or videos. This raises concerns, as individuals could exploit such technology with malicious intent, aiming to spread misinformation or tarnish reputations.